Calculate the potential energy of a charge +6nC placed at 0.07m from a charge +76 x10-6C. If the same charge(+6nC) is now at 0.07m from a charge -88 x10-6C, find its new potential energy.   



1
Expert's answer
2021-10-26T17:21:48-0400

Now we know that

Potential energy

"U=\frac{kq_1 q_2}{r}"


"U=\frac{9\times10^9\times6\times10^{-9}\times76\times10^{-6}}{0.07}=0.05862J"

"U=-\frac{9\times10^9\times6\times10^{-9}\times88\times10^{-6}}{0.07}=-0.0678J"